# Across Text and Vision [Version 0.1]

### v1

@@ 3.2 Micro-Level CoT: Subtask Execution
Once a subtask is assigned by the Macro-Level planner, the Micro-Level CoT is responsible for its execution through iterative, feedback-driven reasoning. Since the overall reliability of the system depends on the consistency of each subtask’s output, the core objective of the Micro-Level CoT is to ensure the generation of stable and high-quality results.
To support this, we introduce a Self-Check (Self-Reflection) mechanism to enhance both robustness and adaptability for subtask execution. Specifically, after attempting to complete a subtask, the model evaluates the quality of its output and determines whether revision is necessary. If logical inconsistencies or cross-modal mismatches are detected, the model revises its output and re-evaluates the result in a closed-loop feedback cycle. This process continues until the model deems the output satisfactory, thereby concluding the subtask execution. The self-check mechanism ensures that generated outputs are aligned with the model’s internal reasoning and multi-modal understanding, as illustrated in Figure 2.
Furthermore, to reduce the complexity of modeling the self-reflection mechanism, we formulate this reasoning process as a Markov Decision Process (MDP) [29], where each self-reflection step depends only on the result of the previous step and the given subtask instruction. As illustrated in Figure 4 (a), we define the current reasoning result (a multi-modal pair) as MDP node $t$, and the revised output after self-evaluation and editing as MDP node $t+1$. We assume that the transition from node $t$ to $t+1$ is solely determined by the previous result and the fixed subtask instruction, which also aligns with human behavior during self-check. To enforce this locality, we introduce an micro masked-attention scheme, which restricts attention to short-range state transitions, as illustrated in Figure 3 (b). This design not only simplifies the learning objective but also enhances training stability and computational efficiency.
Formally, as illustrated in Figure 4, each element in our MDP step $(s_{t},a_{t},s_{t+1},r_{t+1})$ is defined as follows:
  • State $s_{t}$: The current reasoning state, consisting of the output from the previous step, including both textual and visual content;
  • Action $a_{t}$: A hybrid operation that combines textual editing prompt generation with corresponding image editing;
  • Next state $s_{t+1}$: The updated state, encompassing both the edited image and an aligned textual summary;
  • Reward $r_{t+1}$: A textual judgment measuring alignment between the next state’s result and the subtask objective.
With this formulation, the learning objectives of the Micro-Level CoT are formulated into two core competencies: (1) subtask completion, which involves both image editing and multi-modal understanding; and (2) MDP modeling, including hybrid action generation, next-state prediction, and reward estimation, highlighted in pink in Figure 4 (b).
@@ 3.3 Training Paradigm
The training of Uni-CoT is conducted in two sequential stages: Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL).
@@
As described in Sec. 3.1 and Sec. 3.2, our supervised fine-tuning stage comprises two complementary components: (1) interleaved multi-modal supervision for the Macro-Level CoT branch, and (2) multi-task learning for the Micro-Level CoT branch.
For the Macro-Level CoT, we follow BAGEL [7] to adopt a combined loss formulation that supervises both text and image generation. Specifically, we apply cross-entropy (CE) loss for textual output and mean squared error (MSE) loss for image generation:
@@
where $\lambda$ is a balancing coefficient that controls the relative importance of textual and visual losses.
For the Micro-Level CoT, subtask completion is supervised using interleaved multi-modal data, similarly adopting CE and MSE losses. Additionally, to support MDP-based modeling of the self-reflective reasoning process, we decompose learning into four auxiliary objectives and optimize model to learn them respectively as shown table. 1.
Table 1: Details of auxiliary tasks for MDP modeling.
Objective Data Structure Loss
Text Action Generation [System Prompt, Subtask Prompt, Current Image and Text, "Editing Prompt"] Cross-Entropy Loss ($\mathcal{L}_{\text{CE}}^{\text{text}}$)
Image Action Generation [System Prompt, Subtask Prompt, Current Image and Text, Editing Prompt, "Edited Image"] MSE Loss ($\mathcal{L}_{\text{MSE}}^{\text{image}}$)
Next-State Prediction [System Prompt, Subtask Prompt, Edited Image, "Image Analysis"] Cross-Entropy Loss ($\mathcal{L}_{\text{CE}}^{\text{text}}$)
Reward Estimation [System Prompt, Subtask Prompt, Edited Image, Image Analysis, "Evaluation Data"] Cross-Entropy Loss ($\mathcal{L}_{\text{CE}}^{\text{text}}$)
This multi-task supervision framework enables the Micro-Level CoT to learn both the mechanics of action execution and the reflective evaluation of outcomes within an MDP framework.
@@
To further enhance reasoning robustness and adaptability, we adopt reinforcement learning (RL) to optimize both Macro-Level and Micro-Level reasoning behaviors. Rewards are designed to reflect subtask completion quality, multi-modal consistency, and overall task success.
In this work, we employ a simplified yet effective RL strategy, Direct Preference Optimization (DPO) [30, 37], to align model outputs with human-preferred reasoning trajectories. We decouple the DPO training into two preference modeling stages:
  • Textual Preference Learning: Encourages the model to prefer coherent and correct reasoning paths over suboptimal or inconsistent alternatives, using pairwise preference annotations.
  • Visual Preference Learning: Guides image editing behavior by optimizing toward preferred visual outputs, based on human or proxy feedback that captures semantic correctness, visual fidelity, and alignment with the instruction.
@@ 4.1 Implementation Details
Dataset Curation Process.
Figure 6 illustrates the data pipeline we used for data curation. We collect interleaved text-image data for Macro-level and Micro-level reasoning paradigm respectively. We prepare text-to-image generation prompts from multiple datasets as seed prompts for prompt expansion.
For Macro-level reasoning data, we then enhance the prompts in the following two aspects: (1) We first check if the given prompts entail domain expertise knowledge or common-sense reasoning. If such reasoning or logical induction exist in prompts, we rewrite the prompts to explicitly state the result of reasoning deduction. For example, if the original prompt describes ”a melting ice cream cone in desert sun”, then the deduced rewritten prompt should elaborate on the object states under given condition, i.e. ”dripping, puddle on hot sand” (2) We then enrich the generation prompts via adding auxiliary visual detail or attributes to concretely illustrate any abstract concepts or description, especially those regarding art styles, vibes, environment, etc. After prompt enhancement, we employ models such as GPT-4o or Qwen-plus to decompose the enhanced prompts into 2-3 subtasks such that complex image generation goal inferred by the enhanced prompts are broken into simpler and logically coherent sequential or parallel sub-goal. We then utilize models capable of image generation such as BAGEL-Think or GPT-4o, to perform image generation and editing following the subtask instruction. We also employ VLM models (GPT-4o) to evaluate on the results of subtask excecution and generate corresponding subtask refinement instruction. The textual detail of subtask planning, subtask evaluation and refinement, and the intermediate images generated in all subtask steps, are all collected as interleaved Macro-level data.
Figure 6: Data curation pipeline for hierarchical reasoning. We collect T2I prompts from various datasets, then for (a)Macro-level reasoning data pipeline: prompts are first enhanced via GPT-4o or Qwen3, then decomposed into several sequential or parallel subtasks. Then GPT-4o as well as Bagel-Think are used for subtask image generation, evaluation and refinement.(b)Micro-level reasoning data pipeline: we use Bagel-Think model to generate preliminary images based on collected T2I prompts. We then perform several rounds of self-reflection, using GPT-4o to first evaluate on current images and generate refinement instruction, then generate edited images conditioned on refinement instruction.
@@
Owing to our decomposition of interleaved image-text CoT via MDP modeling and high-quality data construction, the training of Uni-CoT is quite efficient, as all experiments can be accomplished on 8 NVIDIA A100 GPUs. Following common practices, we utilize FlashAttention, Fully Sharded Data Parallel (FSDP), and mixed-precision training for better computation efficiency. Throughout the training process, all model parameters are optimized using Adam optimizer with a constant learning rate of 2e-5. Additionally, we employ a linear warmup learning rate schedule, increasing the learning rate from zero over the first 200 steps. During each training step, we combine all data used for understanding and generation expert training, randomly sample from the combined dataset, and pack the samples into sequences with a target length of 32,768 tokens.
Table 2: Quantitative evaluation results on WISE [26]. Notably, the evaluation results are averaged over five independent runs to ensure statistical robustness. As can be drawn, Uni-CoT achieves the best performance among its open-source alternatives.
@@ 4.2 Experimental Setup
In this work, we primarily focus on image generation tasks within the Uni-CoT framework. Specifically, we evaluate Uni-CoT on two reasoning-centric generation tasks: reasoning-driven image generation and reasoning-driven image editing.
For the image generation task, following [7], we conduct experiments on a widely adopted benchmark, WISE [26], which are designed to assess the model’s ability to generate coherent and faithful visual outputs conditioned on complex reasoning prompts. For the image editing task, we benchmark Uni-CoT on RISE [51] and KRIS [40], both of which target visual reasoning and editing under complex instructions. RISE [51] introduces the first benchmark for reasoning-informed visual editing, covering four key reasoning types: temporal, causal, spatial, and logical. In contrast, KRIS [40] is a diagnostic benchmark that categorizes editing tasks into three foundational knowledge types: factual, conceptual, and procedural.
4.3 Results for Reasoning-based Image Generation
@@
As shown in Figure 5, Uni-CoT is capable of correcting semantically inaccurate generations via multi-round self-check. For prompts requires complex reasoning over commonsense and their visual appearance, the initial outputs of Uni-CoT may appear plausible but deviate from the intended semantics. However, by employing a multi-round self-check mechanism, Uni-CoT can iteratively re-evaluate and refine its generated outputs, ensuring that the final visual state in the multimodal reasoning trajectory achieves better alignment with the prompt and improved visual coherence.
Table 3: Quantitative comparisons on KRIS [40]. Uni-CoT achieves the top-1 performance among open-source models on the KRIS benchmark and even surpasses the commercial model Gemini 2.0 by 5.59 points in overall score,validating the robust and competitive image editing abilities owned by Uni-CoT.
@@
Doubao (ByteDance) [2] 70.92 59.17 40.58 63.30 65.50 61.19 62.23 47.75 60.58 54.17 60.70
BAGEL (ByteDance) [7] 64.27 62.42 42.45 60.26 55.40 56.01 55.86 52.54 50.56 51.69 56.21
BAGEL-Think (ByteDance) [7] 67.42 68.33 58.67 66.18 63.55 61.40 61.92 48.12 50.22 49.02 60.18
@@
Table 4: Quantitative comparisons on RISE [51]. Uni-CoT achieves competitive results compared to its baselines and show comparable performance with Gemini-2.0.
Model Overall Performance (%) Evaluation Sub-dimensions
Temporal Causal Spatial Logical Overall Instruction Appearance Visual
Reasoning Consistency Plausibility
GPT-4o (OpenAI) [18] 34.1 32.2 37.0 10.6 28.9 62.8 80.2 94.9
Gemini-2.0-Flash-exp (Google) [19] 8.2 15.5 23.0 4.7 13.3 48.9 68.2 82.7
Gemini-2.0-Flash-pre (Google) [19] 10.6 13.3 11.0 2.3 9.4 49.9 68.4 84.9
Uni-CoT 8.2 18.9 20.0 1.2 12.5 47.4 69.4 84.4
BAGEL-Think (ByteDance) [7] 5.9 17.8 21.0 1.2 11.9 45.9 73.8 80.1
BAGEL (ByteDance) [7] 2.4 5.6 14.0 1.2 6.1 36.5 53.5 73.0
@@
4.4 Results for Reasoning-based Image Editing
@@
We report KRIS and RISE benchmark results in Table 3 and Table 4. On KRIS benchmark, Uni-CoT outperforms all open-source baselines across perception, conceptual, and procedural categories. Remarkably, it also surpasses the closed-source Gemini 2.0 in overall score, highlighting its robust and interpretable editing capabilities under complex reasoning instructions. On RISE benchmark, Uni-CoT demonstrate comparable performance with Gemini 2.0 with respect to both overall performance on the four reasoning categories and the sub-dimension evaluation metrics of instruction reasoning, appearance consistency and visual plausibility.
@@
Figure 7 showcases step-by-step editing results on challenging instructions. Uni-CoT demonstrates strong fidelity to the original context while making precise visual modifications. As exhibited in this figure, our method achieves high prompt consistency, significant spatial accuracy, and plausible visual transitions—highlighting the effectiveness and interpretability of our CoT-guided editing strategy.

### v2

@@ 3.2 Micro-Level CoT: Subtask Execution
Once a subtask is assigned by the Macro-Level planner, the Micro-Level CoT is responsible for its execution. Since the overall reliability of a system hinges on the consistency and coherence of outputs across subtasks, the core objective of the Micro-Level CoT is to generate high-reliable and high-quality results for the assigned subtask.
@@
To support this, we introduce a Self-Reflection mechanism to enhance both robustness and adaptability for subtask execution. Specifically, as illustrated in Figure 2, after attempting to complete a subtask, the model evaluates the quality of its output and determines whether revision is necessary. If logical inconsistencies or cross-modal mismatches are detected, the model revises its output and re-evaluates the result in a closed-loop feedback cycle. This process continues until the model deems the output satisfactory, thereby concluding the subtask execution. This mechanism ensures that generated outputs are aligned with the model’s internal reasoning and multi-modal understanding.
To further reduce the complexity of modeling the self-reflection mechanism, we treat this reasoning process as a Markov Decision Process (MDP) [51], where each neighboring reasoning and self-reflection sequence pair depends only on the outcome of the previous step and the given subtask instruction. As shown in Figure 3 (a), we define the current reasoning result (a multi-modal pair) as MDP node $t$, with the revised output after self-evaluation and editing as MDP node $t+1$. We assume that the transition from node $t$ to $t+1$ is solely governed by the node $t$ and the fixed subtask instruction, mirroring human behavior during self-reflection, where attention is focused on assessing the current state. To enforce this locality, we introduce a micro attention mask, which limits attention to short-range state transitions, as illustrated in Figure 2. This approach simplifies the learning objective while enhancing both training stability and computational efficiency.
Formally, as illustrated in Figure 3 (b), each element in our MDP step $(s_{t},a_{t},s_{t+1},r_{t+1})$ is defined as follows: State $s_{t}$: The current state that consists of the previous step’s results, including both textual and visual content; Action $a_{t}$: A hybrid operation that consists of textual editing prompt generation $a^{text}_{t}$ and corresponding image editing $a^{image}_{t}$; Next state $s_{t+1}$: The updated state, encompassing both the edited image and an aligned textual summary; Reward $r_{t+1}$: A textual judgment measuring alignment between the next state’s result and the subtask objective. With this formulation, the learning objectives of the Micro-Level CoT are formulated into two core competencies: (1) subtask completion, which involves both image editing and multi-modal understanding; and (2) MDP modeling, including hybrid action generation, next-state prediction, and reward estimation, highlighted in pink in Figure 3 (b).
@@ 3.3 Training Paradigm
The Uni-CoT training is conducted in two sequential stages: Supervised Fine-Tuning and Reinforcement Learning.
@@
As detailed in above sections, SFT comprises two complementary components: Macro-Level CoT learning and Micro-Level CoT learning.
For the Macro-Level CoT, we adopt a joint loss that supervises both text and image generation for planning and final synthesis. Specifically, following Eq. 4, we apply cross-entropy (CE) loss for textual output and mean squared error (MSE) loss for image generation:
@@
where ${\lambda}_{\text{CE}}$ is a balancing coefficient that controls the relative importance of textual and visual losses.
For the Micro-Level CoT, subtask completion is supervised using interleaved multi-modal data, similarly adopting CE and MSE losses for model learning. Additionally, to learn MDP-based self-reflective reasoning process, we decompose learning into four auxiliary objectives: text action $a^{text}_{t}$ generation, image action generation, next-state prediction and reward estimation. More details and data structures are are provided in Appendix Sec. B.
@@
To further enhance reasoning robustness and adaptability, we adopt reinforcement learning (RL) to optimize both Macro-Level and Micro-Level reasoning behaviors. Rewards are designed to reflect subtask completion quality and overall task success.
In this work, we employ a simplified yet effective RL strategy, Direct Preference Optimization (DPO) [53, 63], to align model outputs with human-preferred reasoning trajectories. We decouple the DPO training into two stages:
  • Textual Preference Learning: Encourages the model to prefer coherent and correct reasoning paths over suboptimal or inconsistent alternatives via pairwise annotations.
  • Visual Preference Learning: Guides image editing behavior by optimizing toward preferred visual outputs, based on human or proxy feedback that captures semantic correctness and visual fidelity.
@@ 4.1 Implementation Details
Dataset.
We construct a multi-modal reasoning dataset to support both macro- and micro-level CoT learning for our Uni-CoT model. At the macro level, seed prompts are augmented with explicit logical deductions and enriched visual details, then decomposed into 2–3 subtasks using large language models. Each subtask yields interleaved text–image traces comprising planning, evaluation, refinement instructions, and intermediate generations. At the micro level, preliminary images are generated and refined through iterative self-reflection loops, where models repeatedly evaluate, adjust, and edit outputs to capture fine-grained reasoning and editing trajectories.
Through this pipeline, our dataset comprises approximately 11K interleaved text–image pairs for macro-level planning traces and 20K pairs for micro-level refinement loops, spanning both text and image modalities. We also crop several data from [4, 67, 8, 80] to enhance the generalization of our model across various scenarios. Additional details on dataset construction and statistics are provided in Appendix Sec. C.
@@
Owing to our decomposition of interleaved image-text CoT via MDP modeling and high-quality data construction, the training of Uni-CoT is quite efficient, as all experiments can be accomplished on 8 NVIDIA A100 GPUs. Following common practices, we utilize FlashAttention, Fully Sharded Data Parallel (FSDP), and mixed-precision training for better computation efficiency. Throughout the training process, all the parameters in the unified model are optimized using Adam optimizer with a constant learning rate of 2e-5. Additionally, we employ a linear warmup learning rate schedule, increasing the learning rate from zero over the first 200 steps. During each training step, we combine all data used for understanding and generation expert training, randomly sample from the combined dataset, and pack the samples into sequences with a target length of 32,768 tokens.
Table 1: Quantitative evaluation results on GenEval [24]. Notably, ‘Gen. Only’ stands for an image generation model, and ‘Unified’ denotes a model that has both understanding and generation capabilities. $\dagger$} refers to the best preformance of Bagel reproduced by ourselves. Uni-CoT Init is our initial results without multi-modal reasoning.
[…]
Gen. Only PixArt-$\alpha$ [9] 0.98 0.50 0.44 0.80 0.08 0.07 0.48
DALL-E $2$ [54] 0.94 0.66 0.49 0.77 0.10 0.19 0.52
Emu33-Gen  [66] 0.98 0.71 0.34 0.81 0.17 0.21 0.54
SDXL [50] 0.98 0.74 0.39 0.85 0.15 0.23 0.55
DALL-E $3$ [3] 0.96 0.87 0.47 0.83 0.43 0.45 0.67
SD3-Medium [18] 0.99 0.94 0.72 0.89 0.33 0.60 0.74
FLUX.1-dev [34] 0.98 0.93 0.75 0.93 0.68 0.65 0.82
Unified LWM [37] 0.93 0.41 0.46 0.79 0.09 0.15 0.47
SEED-X [23] 0.97 0.58 0.26 0.80 0.19 0.14 0.49
TokenFlow-XL [52] 0.95 0.60 0.41 0.81 0.16 0.24 0.55
ILLUME [64] 0.99 0.86 0.45 0.71 0.39 0.28 0.61
Emu33-Gen [66] 0.99 0.81 0.42 0.80 0.49 0.45 0.66
Show-o [74] 0.98 0.80 0.66 0.84 0.31 0.50 0.68
Janus-Pro-7B [10] 0.99 0.89 0.59 0.90 0.79 0.66 0.80
MetaQuery-XL [48] - - - - - - 0.80
BAGEL [15] 0.99 0.92 0.78 0.87 0.53 0.64 0.79
Uni-CoT 0.99 0.95 0.82 0.89 0.60 0.72 0.83
Uni-CoT Init 0.99 0.95 0.82 0.90 0.55 0.69 0.81
Uni-CoT 0.99 0.96 0.84 0.92 0.57 0.71 0.83
Table 2: Quantitative evaluation results on WISE [46]. Notably, the evaluation results are averaged over five independent runs to ensure statistical robustness. As can be drawn, Uni-CoT achieves the best performance among its open-source alternatives. Uni-CoT Init is our initial results without multi-modal reasoning.
@@ 4.2 Experimental Setup
In this work, we focus on the Uni-CoT on two generation tasks in main paper: image generation and image editing.
For the image generation task, following [15], we conduct experiments on two widely adopted benchmarks: GenEval [24] and WISE [46]. GenEval [24] serves as a general benchmark for assessing object-focused text-to-image alignment, while WISE [46] is a reasoning-driven benchmark designed to evaluate a model’s ability to generate faithful outputs from abstract, reasoning-intensive prompts.
For the image editing task, we benchmark Uni-CoT on GEdit-Bench [38], RISE [85], and KRIS [71]. GEdit-Bench [38] is a general benchmark that provides a diverse set of real-world editing tasks. In contrast, RISE [85] focuses on reasoning-informed editing across temporal, causal, spatial, and logical dimensions, while KRIS [71] serves as a diagnostic benchmark categorizing editing tasks into factual, conceptual, and procedural knowledge types.
4.3 Experimental Setup
@@
Table 1 presents the evaluation results of Uni-CoT and other baselines on GenEval [24], which benchmarks basic image generation capabilities. Among the compared models, Uni-CoT outperforms its base model Bagel. We attribute this improvement primarily to the self-reflection mechanism, which effectively corrects errors in the initial generations.
@@
As shown in Figure 4, Uni-CoT is capable of correcting semantically inaccurate generations via multi-round self-reflection. For prompts requires complex reasoning over commonsense and their visual appearance, the initial outputs of Uni-CoT may appear plausible but deviate from the intended semantics. However, by employing a multi-round self-reflection mechanism, Uni-CoT can iteratively re-evaluate and refine its generated outputs, ensuring that the final visual state in the multimodal reasoning trajectory achieves better alignment with the prompt and improved visual coherence.
Table 3: Quantitative comparisons on GEdit-Bench [38]. All metrics are reported as higher-is-better ($\uparrow$).
[…]
G_SC G_PQ G_O G_SC G_PQ G_O
Private Gemini 2.0 [33] 6.73 6.61 6.32 5.43 6.78 5.36
GPT-4o [31] 7.85 7.62 7.53 7.67 7.56 7.30
Open-source Instruct-Pix2Pix [6] 3.58 5.49 3.68 - - -
MagicBrush [83] 4.68 5.66 4.52 - - -
AnyEdit [81] 3.18 5.82 3.21 - - -
OmniGen [73] 5.96 5.89 5.06 - - -
Step1X-Edit [39] 7.09 6.76 6.70 7.20 6.87 6.86
BAGEL [15] 7.36 6.83 6.52 7.34 6.85 6.50
Uni-CoT 7.91 6.24 6.74 8.01 6.30 6.87
Table 4: Quantitative comparisons on KRIS [71]. Uni-CoT achieves the top-1 performance among open-source models on the KRIS benchmark and even surpasses the commercial model Gemini 2.0 by 5.59 points in overall score.
@@
Doubao [7] 70.92 59.17 40.58 63.30 65.50 61.19 62.23 47.75 60.58 54.17 60.70
BAGEL [15] 64.27 62.42 42.45 60.26 55.40 56.01 55.86 52.54 50.56 51.69 56.21
BAGEL-Think [15] 67.42 68.33 58.67 66.18 63.55 61.40 61.92 48.12 50.22 49.02 60.18
@@
Table 5: Quantitative comparisons on RISE [85]. Uni-CoT superpass Bagel and show comparable performance with Gemini-2.0.
Model Temporal Causal Spatial Logical Overall
Gemini-2.0 [33] 8.2 15.5 23.0 4.7 13.3
BAGEL-Think [15] 5.9 17.8 21.0 1.2 11.9
BAGEL [15] 2.4 5.6 14.0 1.2 6.1
Uni-CoT 8.2 18.9 20.0 1.2 12.5
GPT-4o [31] 34.1 32.2 37.0 10.6 28.9
@@
4.4 Results for Image Editing
@@
First, we evaluate the basic editing capability of Uni-CoT on GEdit-Bench [38]. As shown in Table 3, our model achieves competitive performance compared with other approaches.
Furthermore, we report KRIS and RISE benchmark results in Table 4 and Table 5. Surprisingly, on KRIS benchmark, Uni-CoT outperforms all open-source baselines across perception, conceptual, and procedural categories. Remarkably, it also surpasses the closed-source Gemini 2.0 in overall score, highlighting its robust and interpretable editing capabilities under complex reasoning instructions. On RISE benchmark, Uni-CoT demonstrate comparable performance with Gemini 2.0 with respect to both overall performance on the four reasoning categories and the sub-dimension evaluation metrics of instruction reasoning, appearance consistency and visual plausibility.
Figure 6: Qualitative Results for Nano-Banana like Generation.
@@
Figure 5 showcases editing results on challenging instructions. Uni-CoT demonstrates strong fidelity to the original context while making precise visual modifications. As exhibited in this figure, our method achieves high prompt consistency, significant spatial accuracy, and plausible visual transitions, highlighting the effectiveness and interpretability of our CoT-guided editing strategy.
4.5 Nano-banana like Generation
Recently, Google introduced Nano-Banana [44], a unified image generator that achieves results once considered unattainable, such as synthesizing realistic landscapes from satellite imagery or isohypse line maps (top of Figure 6).
We hypothesize that this ability stems from an inherent reasoning-driven generation mechanism. Concretely, the transformation from isohypse maps to landscape images can be decomposed into three steps: (1) 2d-to-3d: interpolating the 2D isohypse map into a 3D terrain; (2) 3d-crop: focusing on a camera view from the 3D representation; and (3) 3d-to-real: rendering the cropped view into a realistic landscape.
To validate this hypothesis, we construct a dataset of 3K multi-modal Chain-of-Thought (CoT) geography samples following this decomposition and refine our model on it. The model converges efficiently and exhibits strong performance on this test set (bottom of Figure 6). In contrast, direct fine-tuning on raw isohypse–landscape pairs results in unstable training and incoherent generations. More results are shown in Figure S3.